OpenStudy (anonymous):

Determine whether the polynomial is a perfect square and if it is, factor it. n^2 − 9n + 9

OpenStudy (atlas):

It doesn't look to me as a perfect square

OpenStudy (atlas):

Perfect squares are of the form - a^2 - 2ab +b^2

OpenStudy (atlas):

Here a =n, b=3 but the middle term is -9n if it would have been -6n, the polynomial would have been a perfect square

OpenStudy (anonymous):

no, it's not perfect square

OpenStudy (whpalmer4):

You can use the determinant \(\Delta = b^2-4ac\) to determine if you have a perfect square in a quadratic of the form \(ax^2+bx+c=0\) If \(\Delta= 0\) you have a perfect square. Here we have \(a =1, b = -9, c = 9\) and \[\Delta = (-9)^2-4(1)(9) = 81-36 = 45\]Therefore it is not a perfect square.

OpenStudy (whpalmer4):

Sorry, discriminant, not determinant...